OpenStudy (anonymous):

y=2x-9 -4x-7y=9 how do i get the second equation to be y=mx+b

hero (hero):

Why are you trying to get the second equation to be y = mx + b? \(-4x - 7y = 9\) \(7y = -4x - 9\) \(7y = -(4x + 9)\) \(y = -\dfrac{4x + 9}{7}\)